• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 11:35:02 on April26,2007

    Entry number 200132

    keyword=Changed to Q2=55m-2

    We change to Q2=55m-2, but the collimation system did not work.
    We are now waiting for beam as MCC tries to fix a 60 Hz problem on the Hall B beam.